How many sides does a regular polygon have if the sum of its interior angle measures is 7380 degree?

Answer by Alan3354(69443)

How many sides does a regular polygon have if the sum of its interior angle measures is 7380 degree?
---------
The sum of the angles is (n-2)*180
n-2 = 7380/180 = 41
n = 43 sides